OpenStudy (anonymous):

If y = cot^-1((cosx)^1/2) - tan^-1((cosx)^1/2), prove that siny = tan^2(x/2).

OpenStudy (anonymous):

If \[y = \cot ^{-1}\sqrt{cosx} - \tan ^{-1}\sqrt{cosx}\] prove that \[siny = \tan ^{2}\frac{ x }{ 2 }\]